OpenStudy (anonymous):

An aqueous solution of 6.3g oxalic acid dehydrate is made upto 250 ml., The volume of 0.1 N NaOH required to completely neutralize 10 mL of this solution is

OpenStudy (anonymous):

V=0,0559 dm3 or 55,9 ~ 56 cm3

OpenStudy (anonymous):

H2C2O4 + 2NaOH --> Na2C2O4 + 2H2O n (H2C2O4)=m/M=0,06997 mol c=n/V=0,27988 M in 10cm3: n=0,27988M*0,01 dm3=2,7988*10^-3 mol n(NaOH)=2*n(H2C2O4) = 5,5976*10^-3 mol V=n/c = 5,5976*10^-3 mol/0,1M=0,055976 dm3 = 56 cm3

OpenStudy (anonymous):

thanks